OpenStudy (anonymous):

Solve for x: 2 over 3 (x - 2) = 4x. negative 2 over 5 negative 5 over 2 2 over 5 5

OpenStudy (luigi0210):

Do you know how to write the equation?

OpenStudy (anonymous):

i think so

OpenStudy (anonymous):

Solve for x: 2 over 3 (x - 2) = 4x.

OpenStudy (luigi0210):

Is it this: \[\LARGE \frac{2}{3}(x-2)=4x\] or \[\LARGE \frac{2}{3(x-2)}=4x\]

OpenStudy (anonymous):

first one

OpenStudy (luigi0210):

Then what you need to do is first distribute that 2/3. What do you get when you do that?

OpenStudy (anonymous):

so im distributing 2/3 , witch is multiplying it by all the other numbers in the equation ?

OpenStudy (luigi0210):

That are in the parenthesis.

OpenStudy (anonymous):

idk how to multiply the fraction into a x, and idk how to multiply a fraction into a number

OpenStudy (luigi0210):

I'll start you off, when you distribute you get this: \[\LARGE \frac{2}{3}x-\frac{2*2}{3}=4x\] To get rid of the fraction, just multiply everything by 3 to get: \[\LARGE 2x-(2*2)=12x\] that 2*2 is just 4 so: \[\LARGE 2x-4=12x\] Can you finish it?

OpenStudy (anonymous):

really Ⓐ Luigi0210 and u said everyone isnt so nice like meh

OpenStudy (anonymous):

no, do i divide the 4 into the 12x or the 2x?

OpenStudy (katie0146722):

1 over 2 x + 2 = 22 40

OpenStudy (luigi0210):

@asiarose14 Hmm? @sunshinee no, isolate the variables from the coefficient.

OpenStudy (anonymous):

im not sure what that means i thought a coefficient had something to do with a fraction

OpenStudy (luigi0210):

Okay, so you have \[\LARGE 2x-4=12x\] What I mean is get the x's on one side and numbers on the other. So subtract 2x from both sides. \[\LARGE 2x-4-2x=12x-2x\] Simplify: \[\LARGE -4=10x\] Now you can divide and get x.

OpenStudy (luigi0210):

@asiarose14 if you need help I'm more than happy to assist, just don't always expect a direct answer.

OpenStudy (anonymous):

is the answer negative 2 over 5 ? @Luigi0210

OpenStudy (luigi0210):

Yup :)